# Evolutionary Policy Optimization

### v2

@@ 5.2 Results and Analysis
In manipulation tasks, while methods like PPO, SAPG, and PBT perform reasonably on simpler tasks (e.g., single-arm regrasping and throwing), EPO still achieves higher success rates. For more complex tasks requiring two-arm coordination, baselines fail to learn meaningful behavior even after $1\times 10^{10}$ steps, whereas EPO achieves strong performance. For example, EPO reaches 35.8 episode successes in two-arm reorientation using $1\times 10^{10}$ steps, surpassing SAPG’s 28.58 after $5\times 10^{10}$ steps in the original SAPG paper [41], highlighting EPO’s superior sample efficiency. A similar trend holds in locomotion. On flat-ground ANYmal walking, baselines match EPO, but in more complex parkour tasks requiring rich exploration, EPO significantly outperforms them. In the DeepMind Control Suite (Humanoid), all methods perform reasonably, yet EPO still yields marginally higher returns.
Comparing SAPG (5th column) with SAPG 6 (6th column) reveals that hybrid-policy-only updates do not benefit from large-scale simulation and may even degrade performance, as increasing the number of agents introduces a higher proportion of poorly performing policies. In contrast, comparing SAPG (5th column) with EPO (last column) highlights the critical role of the genetic algorithm (Table 1), especially in more challenging tasks like two-arm manipulation. By filtering out underperforming agents and learning only from high-quality data, the GA ensures stable and effective training. This demonstrates that the genetic algorithm is not a minor enhancement, but a fundamental component of the hybrid-policy update.
@@
Compared to EvoRL baselines, EPO demonstrates superior learning efficiency and asymptotic performance. While PBT performs well on simpler tasks, its effectiveness diminishes in more complex scenarios. This is largely due to PBT’s requirement to train each policy nearly to convergence before selecting optimal hyperparameters, along with its inability to leverage data from past agents—resulting in inefficient use of experience. Consequently, PBT does not benefit from larger population sizes, as evidenced by the similar performance between PBT and PBT 8. In contrast, CEM-RL relies heavily on off-policy updates, which becomes a limitation in difficult tasks where most agents perform poorly. This leads to the accumulation of low-quality data and ultimately degrades asymptotic performance across a wide range of tasks. EPO overcomes these limitations by employing a hybrid-policy update mechanism while retaining the genetic algorithm for agent selection, particularly in complex environments.
@@ 5.3 Ablations
In this ablation study, we investigate the impact of population size while keeping the total computational budget fixed, using the single-arm reorientation task—where the agent must grasp an object and reorient it to a target 6-DoF pose—as the testbed. Unless otherwise specified, all subsequent analyses in this section are conducted within this environment. Specifically, we fix the number of environments to $N=24{,}576$ and vary the population size from $K=8$ to $K=64$ (Figure 4(a)). The results indicate that our algorithm benefits from a larger population. While this may seem intuitive, it is not a guaranteed outcome—larger populations introduce greater behavioral diversity, which can also increase the prevalence of poorly performing agents and result in noisier data. To address this, EPO applies Genetic Algorithms (GA) at the latent space level to periodically eliminate underperforming agents, ensuring that only high-quality data contributes to learning. Moreover, since all agents share a common actor-critic network, we maintain controlled diversity while avoiding unbounded policy divergence—an essential property for stable and scalable training.
We also analyzed the impact of different crossover strategies beyond simple averaging. Specifically, we experimented with two alternative approaches—uniform crossover and fitness-weighted averaging. In uniform crossover, each element of the latent vector is randomly selected from one of the two parents with equal probability. In the fitness-weighted average approach, the parents’ latent vectors are combined proportionally to their fitness scores. Our method achieved a 36.7% success rate, compared to 32.6% for uniform crossover and 37.0% for fitness-weighted averaging, indicating that performance is relatively robust to the choice of crossover strategy. And for simplicity, we keep the averaging strategy.
@@ 5.4 Scaling Law
It is worth noting that our method focuses on scaling with training data, which is orthogonal to efforts that scale performance through larger neural network architectures. This makes it naturally compatible with existing regularization techniques designed for training large networks [32, 23]. Additionally, the shared actor-critic architecture enhances memory efficiency and inherently supports scaling to larger network sizes.
@@ Appendix A Computational Cost
We also evaluate the computational cost of EPO compared with other baselines that involve a population-based component (SAPG, PBT, CEM-RL). Since EPO requires calculating fitness score of all agents and determining whether iteration is necessary, it slightly increases the computational cost over PBT and SAPG. However, simply increasing the number of agents in EPO does not have a significant impact on memory usage or inference time. This is because the additional cost mainly comes from adding a small number of latent embeddings, which does not drastically affect the overall resource usage. All results are obtained on task M:SO with 24,576 environments and $95e$ 9 environment steps.
@@ B.2 Hyperparameters and Network
Mini-batch size 44$\cdot$4 ⋅ num_envs
@@
$x$ in Line 13 of Algo.1 num_agents$(K)-2$
@@
Mini-batch size 44$\cdot$4 ⋅ num_envs
@@
$x$ in Line 13 of Algo.1 num_agents$(K)-2$

### v3

@@ 5.2 Results and Analysis
In manipulation tasks, while methods like PPO, SAPG, and PBT perform reasonably on simpler tasks (e.g., single-arm regrasping and throwing), EPO still achieves higher success rates. For more complex tasks requiring two-arm coordination, baselines fail to learn meaningful behavior even after $1\times 10^{10}$ steps, whereas EPO achieves strong performance. For example, EPO reaches 35.8 episode successes in two-arm reorientation using $1\times 10^{10}$ steps, surpassing SAPG’s 28.58 after $5\times 10^{10}$ steps in the original SAPG paper Singla et al. (2024), highlighting EPO’s superior sample efficiency. A similar trend holds in locomotion. On flat-ground ANYmal walking, baselines match EPO, but in more complex parkour tasks requiring rich exploration, EPO significantly outperforms them. In the DeepMind Control Suite (Humanoid), all methods perform reasonably, yet EPO still yields marginally higher returns.
Comparing SAPG 64 (5th column) with SAPG 6 (6th column) reveals that hybrid-policy-only updates do not benefit from large-scale simulation and may even degrade performance, as increasing the number of agents introduces a higher proportion of poorly performing policies. In contrast, comparing SAPG 64 (5th column) with EPO (last column) highlights the critical role of the genetic algorithm (Table 1), especially in more challenging tasks like two-arm manipulation. By filtering out underperforming agents and learning only from high-quality data, the GA ensures stable and effective training. This demonstrates that the genetic algorithm is not a minor enhancement, but a fundamental component of the hybrid-policy update.
@@
Compared to EvoRL baselines (PBT and CEM-RL), EPO demonstrates superior learning efficiency and asymptotic performance. While PBT performs well on simpler tasks, its efficiency decreases on more complex ones. In PBT, each agent learns only from its own experience, and the data collected by eliminated agents is discarded rather than reused. This limitation becomes more pronounced in complex tasks that require significantly more environment steps for training. Consequently, PBT does not benefit from larger population sizes, as evidenced by the similar performance between PBT 64 and PBT 8. In contrast, CEM-RL relies heavily on off-policy updates, which becomes a limitation in difficult tasks where most agents perform poorly. This leads to the accumulation of low-quality data and ultimately degrades asymptotic performance across a wide range of tasks. EPO overcomes these limitations by employing a hybrid-policy update mechanism while retaining the genetic algorithm for agent selection, particularly in complex environments.
@@ 5.3 Ablations
In this ablation study, we investigate the impact of population size while keeping the total computational budget fixed, using the single-arm reorientation task—where the agent must grasp an object and reorient it to a target 6-DoF pose—as the testbed. Unless otherwise specified, all subsequent analyses in this section are conducted within this environment. Specifically, we fix the number of environments to N=24,576N=24{,}576 and vary the population size from $K=8$ to $K=64$ (Figure 5(a)). The results indicate that our algorithm benefits from a larger population. While this may seem intuitive, it is not a guaranteed outcome—larger populations introduce greater behavioral diversity, which can also increase the prevalence of poorly performing agents and result in noisier data. To address this, EPO applies Genetic Algorithms (GA) at the latent space level to periodically eliminate underperforming agents, ensuring that only high-quality data contributes to learning. Moreover, since all agents share a common actor-critic network, we maintain controlled diversity while avoiding unbounded policy divergence—an essential property for stable and scalable training.
We analyzed the impact of different crossover strategies beyond simple averaging, testing two alternatives: uniform crossover and fitness-weighted averaging. In uniform crossover, each element of the latent vector is randomly chosen from one parent with equal probability, while in fitness-weighted averaging, vectors are combined proportionally to fitness scores. Our method achieved a 36.7% success rate, compared to 32.6% for uniform crossover and 37.0% for fitness-weighted averaging, showing robustness to the crossover choice. For simplicity, we keep the averaging strategy.
@@ 5.4 Scaling Law
It is worth noting that our method focuses on scaling with training data, which is orthogonal to efforts that scale performance through larger neural network architectures. This makes it naturally compatible with existing regularization techniques designed for training large networks Nauman et al. (2024); Lee et al. (2024). Additionally, the shared actor-critic architecture enhances memory efficiency and inherently supports scaling to larger network sizes.
@@ 6 Conclusion
Ethics Statement
While EPO does not present direct ethical concerns, RL methods, particularly those that improve large-scale training, could have broader societal implications. Scalable RL has potential applications in autonomous systems, decision-making AI, and robotics, which may impact labor markets, automation policies, and safety regulations. It is crucial to ensure that future applications of such methods align with ethical AI principles, particularly regarding fairness, transparency, and robustness in high-stakes decision-making.
Additionally, as RL is increasingly applied to large-scale language models (LLMs) and automated reasoning systems, scalability improvements may accelerate advancements in AI alignment and human-AI interaction. However, misuse of reinforcement learning in generating persuasive AI or manipulative automated systems remains a potential risk. We encourage the community to apply EPO responsibly and consider ethical safeguards when deploying large-scale RL systems in real-world applications.
Reproducibility Statement
For reproducibility, we have provided an anonymous link to the downloadable source code. Please refer to https://yifansu1301.github.io/EPO/for more details.
@@ Appendix A Computational Cost
We also evaluate the computational cost of EPO compared with other baselines that involve a population-based component (SAPG, PBT, CEM-RL). Since EPO requires calculating fitness score of all agents and determining whether iteration is necessary, it slightly increases the computational cost over PBT and SAPG. However, simply increasing the number of agents in EPO does not have a significant impact on memory usage or inference time. This is because the additional cost mainly comes from adding a small number of latent embeddings, which does not drastically affect the overall resource usage. All results are obtained on task M:SO with 24,576 environments and $5e9$ environment steps.
@@ B.2 Hyperparameters and Network
Mini-batch size $4\cdot$ num_envs
@@
$x$ in Line 13 of Algo.1 num_agents(K)2(K)-2
@@
Mini-batch size $4\cdot$ num_envs
@@
$x$ in Line 13 of Algo.1 num_agents(K)2(K)-2
B.3 The Use of Large Language Models (LLMs)
We use LLMs solely for polishing writing and checking grammar.